# GraspDiffusion: Synthesizing Realistic Whole-body Hand-Object Interaction

### v2

@@ Abstract
Recent generative models can synthesize high-quality images but often fail to generate humans interacting with objects using their hands. This arises mostly from the model’s misunderstanding of such interactions, and the hardships of synthesizing intricate regions of the body. In this paper, we propose GraspDiffusion, a novel generative method that creates realistic scenes of human-object interaction. Given a 3D object mesh, GraspDiffusion first constructs life-like whole-body poses with control over the object’s location relative to the human body. This is achieved by separately leveraging the generative priors for 3D body and hand poses, optimizing them into a joint grasping pose. The resulting pose guides the image synthesis to correctly reflect the intended interaction, allowing the creation of realistic and diverse human-object interaction scenes. We demonstrate that GraspDiffusion can successfully tackle the relatively uninvestigated problem of generating full-bodied human-object interactions while outperforming previous methods.
Figure 1: Given an object mesh and its relative position, GraspDiffusion generates whole body grasping 3D poses, which is subsequently used as guidance for creating human-object interaction scenes. As shown, GraspDiffusion can synthesize images with valid human-object interactions for various types of objects. Note that the bottom-right sample (a green bag) was created from an object image, which was made into a 3D using TripoSR [71], further paving the way for various use cases.
@@ 1 Introduction
The recent advent of diffusion-based generative models [66, 67, 27] has demonstrated significant success in producing high-quality visual content [62, 55, 61, 65, 58]. When trained on large datasets, these models can coherently synthesize images of various subjects corresponding to given textual/visual cues. However, despite their strong performance, generative models struggle to comprehend and visualize everyday hand-object interactions. This limitation hinders their broader adoption in generative model-based content creation pipelines.
@@
While several papers [50, 73, 57] have applied inpainting for hand region refinement, they only focus on situations where the hand is not interacting with other objects, making it impractical for most use cases. In order to generate realistic grasping hands for a given object, the models must understand the semantics and functionalities provided by the object–a concept well known as affordance [18]. While Affordance Diffusion [82] and HOIDiffusion [85] creates an image of a single grasping hand using affordance, these methods only represent the explicit physical contact between a human hand and the object (devoid of any human identity), and fail to convey the spatial / orientational non-contact relationships, making them unsuitable for understanding human affordance. For instance, when using a cell phone, the relationships between the human’s face and torso with respect to the phone should be considered as part of interaction along with the hand touching the phone. This requirement necessitates the development of an pipeline that creates identifiable human-object images where the human’s body is visible, such that the implicit, non-contact relationships are well captured.
In this paper, we present GraspDiffusion, a novel method for generating interaction images with realistic hands and a clear human identity from a single 3D object input. Instead of relying on textual prompts, which tend to be ambiguous in describing complex interactions [54], we first utilize a diffusion model to generate full-bodied grasping poses [56, 63] conditioned on the object and its position. The generated 3D pose parameters are provided as conditions for the next stage to create high-quality image samples. Specifically, we train multiple conditional models [84, 53] and a novel cross-attention modulation scheme [3, 14] to correctly convey the interaction context without harming the diversity of the generated samples. To overcome the lack of a large-scale image-3D pose paired dataset for human object interaction, we also propose a dataset annotation scheme to gather 3D annotations from previous image-based interaction datasets [20, 19, 10, 45, 30].
@@ 2 Related Work
Conditional Image Generation. To provide additional fine-grained, spatial conditions for diffusion models, ControlNet [84] and T2I-Adapter [53] proposed using image-level signals to control the generation process, which includes using 2D human keypoint skeletons [9, 80] for human pose guidance and depth maps for better depth perception. While the improvements in conditional image generation have been significant, synthesizing humans interacting with objects hasn’t reached the same level of improvement.
@@
Several papers [50, 73, 57] proposed to refine the hands of images generated using Controlnet-based inpainting, but does not count as a direct solution to human-object interaction. Others focused on identifying contacts for inpainting a new hand or object for a given scene [82, 38, 52, 79], yet they are limited in camera views and require a prior scene context, making them infeasible in practical scenarios. HOIDiffusion [85] also generates images from synthesized grasps, but is limited to hand-centric views. Compared to HanDiffuser [54], which applied the injection of hand embeddings during image generation to create realistic hands, our pipeline focuses more on the joint synthesis of hand and object, and uses spatial guidance to better direct the generation process towards a 3D scene context.
@@
Grasp Synthesis. Synthesizing a hand grasp consisting of a given object and a hand model is important in understanding human-object interaction, and is a widely studied task in robotics, graphics, and computer vision. While the focus in robotics is to make stable grasps for a given object in simulation / real life [43, 72, 74, 16, 76], in computer vision and graphics the focus is to make plausible grasps that are physically plausible, and generate grasps for either hands [12, 29, 32, 88, 86, 13, 47, 44] or a human [70, 17, 69, 78, 8, 87, 51].
@@
Thanks to the advent of human-object interaction datasets [4, 7, 11, 15, 21, 23, 39, 48, 68], many grasp synthesis methods achieved high performance in generating plausible grasps. Yet most existing datasets have issues with data scalability and variability, especially when it comes to color image-3D paired datasets. While hand-object interaction datasets like DexYCB [11], ARCTIC [15] and HOI4D [48] consist of richly annotated image data, they are rather focused only on the hand and object recorded in an egocentric manner, devoid of any human identity and spatial / orientational non-contact relationships. Although BEHAVE [4] contains both RGB video sequences with 3D annotation, the image quality and motion sensors worn by the subject make it difficult to use as a realistic image dataset. To overcome this issue, we used traditional human-object interaction datasets [20, 19, 10, 45, 30] along with annotation tools [37, 46, 40, 41, 33] to create pseudo-3D annotations for the 2D image.
@@
Building our insight from similar approaches, [70, 87], we focus on leveraging priors from a full-body pose model and a hand-grasping model. Compared to previous approaches [69, 17, 78, 8], instead of generating a motion sequence, we synthesize the pose parameters for the 3D parametric models [56, 63] using a diffusion model.
@@ 3 GraspDiffusion
Fig. 3 illustrates the proposed architecture. Starting with a 3D object mesh and its position within the human-centric coordinate system (originating at the pelvis joint), GraspDiffusion synthesizes realistic images portraying a human interacting with the object, with a significant portion of the human body visible to be considered ”full-body”. In the initial full-body grasping stage, we generate the pose parameters for the human body model [56] interacting with the 3D object mesh (Section. 3.2). In the scene generation stage, we extract geometric structures from the pose parameters to guide the generation of realistic images, leveraging a latent diffusion model [62] along with spatial encoders and a cross-attention modulation scheme (Section. 3.3).
@@ 3.1 Preliminaries
Diffusion Models. Diffusion models [66, 27] are a group of generative models that interpret the data distribution $p(x)$ as a sequential transformation from a tractable prior distribution $p(x_{T})\sim\mathcal{N}(0,I)$. During training, the model uses a forward noise process $q(x_{t}|x_{t-1})$ that gradually adds a small amount of noise to a clean data sample $x_{0}$ towards $p(x_{T})$. At the same time, the model learns a backward noise process $p(x_{t-1}|x_{t})$ implemented as a neural network, which is trained to remove the noise from the before generating samples from $p(x_{T})$. For latent diffusion models [62], the diffusion process is performed in the latent space of a trained autoencoder model [36], guided by a conditional text embedding derived from the CLIP [60] mechanism.
3D Parametric Models. For the hand model, we use the MANO [63] differentiable model, in which we input the full finger articulated hand pose $\theta_{\text{h}}\in\mathbb{R}^{15\times 3}$, wrist translation $t_{\text{h}}\in\mathbb{R}^{3}$ and global orientation $R_{\text{h}}\in\mathbb{R}^{3}$ and get a 3D mesh $\mathcal{M}_{\text{h}}$ with vertices $\mathcal{V}_{\text{h}}$. For the full-body model, we use the SMPL-X [56] differentiable model, in which we input the full-body pose $\theta_{\text{b}}\in\mathbb{R}^{21\times 3}$, the full finger articulated hand pose $\theta_{\text{h}}\in\mathbb{R}^{15\times 3}$ for both hands, the root translation $t_{\text{b}}\in\mathbb{R}^{3}$ and global orientation $R_{\text{b}}\in\mathbb{R}^{3}$ and get a 3D mesh $\mathcal{M}_{\text{body}}$ with vertices $\mathcal{V}_{\text{b}}$.
Figure 4: Full-body grasping pipeline. We separately leverage a hand-grasping model [68] and a body-pose diffusion model, and perform a joint optimization into a full-bodied grasping pose.
@@ 3.2 Full-Body Grasping Pipeline
Building on prior approaches [70, 87], we separately generate hand grasps and body poses in creating a whole-body grasping pose. Specifically, we take a 3D object mesh, its relative location to the human root, and the contacting hand orientation (left or right) as the input, to generate an SMPL-X mesh that grasps the given 3D object with a realistic body pose and hand-object contact.
The input object mesh is used to generate a plausible MANO [63] hand grasp, for which we utilized GrabNet [68], a conditional variational autoencoder (cVAE) that produces hand grasps conditioned on the Basis Point Set (BPS) [59] of the given object. Separately trained for left and right-hand grasps, GrabNet generates MANO parameters $(\theta_{\text{h}},t_{\text{h}},R_{\text{h}})$ that grasps the given object, displaying accurate contact and high generalization for unseen objects.
The object’s relative location $t_{\text{obj}}\in\mathbb{R}^{3}$ and the hand orientation $c_{\text{left}},c_{\text{right}}\in\{0,1\}$ is then used to create a body pose that not only roughly positions its hand in the desired object location, but also reflects the appropriate implicit relationships required for a plausible grasping body pose [69, 34]; whether the head is correctly oriented towards the object, the arms are correctly extended and the torso is leaning towards the object. To achieve this, we utilize a diffusion generative model trained to generate SMPL-X pose parameters $(\theta_{\text{body}},R_{\text{body}})$ conditioned on an object location and whether to use the right/left hand for contact. The loss is defined as
$\mathcal{L}_{DM}=\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,I),t}[||\epsilon-\epsilon_{\theta}(x_{t},t,c)^{2}_{2}||],$ (1)
@@ 3.3 Scene Generation Pipeline
Given the 3D body pose pose, the scene generation pipeline extracts multiple spatial conditions as conditions to a pre-trained [62] model to create consistent images of human-object interaction. Optionally, the pipeline can refine the image focused on the hand-object region to further adjust interaction and correct erroneous details.
@@
To precisely control the human-object image’s generation, we render three spatial conditions from the full-body grasping output. We first render the skeleton $(s^{i})$ of the SMPL-X body, consisting of body and hand joints, to ensure realistic human proportions within the generated image. We also use the joint depth map $(d^{i})$ from the SMPL-X and object model to provide depth information. Lastly, we render the occluded object with ambient lighting $(o^{i})$ to preserve its appearance while relighting it. To apply conditions, we chose the CoAdapter [53] approach, which allows flexibility in handling multiple conditions. For each condition, we separately apply an adapter $\mathcal{F}_{\text{AD}}$ and perform a weighted sum to create feature $\mathbf{F}_{c}$, which can be written as
$\mathbf{F}_{c}=\sum_{k\in\{s,d,o\}}\omega_{k}\mathcal{F}_{\text{AD}}^{k}(k^{i}).\vspace{-0.2cm}$ (3)
During training of the conditioning pipeline, we fix the parameters of the baseline U-Net model and only optimize the conditional adapters, reducing the risk of the model converging to the dataset’s style. This decision allows us to control the image style by applying LoRAs [28] or using finetuned Stable Diffusion models during inference time, making it suitable for real-world application tasks requiring personalized image generation.
$\mathcal{L}_{ADM}=\mathbb{E}_{z,\epsilon\sim\mathcal{N}(0,I),t,\mathbf{F}_{c}}[||\epsilon-\epsilon_{\theta}(z_{t},t,c_{\text{text}},\mathbf{F}_{c})^{2}_{2}||].$ (4)
@@
To address the issue of erroneous interactions, in which interactions may occur from locations other than the intended area (examples on Fig. 8), we introduce a training-free guidance method motivated from prior zero-shot semantic image synthesis techniques  [3, 14]. We first render binary segmentation masks from the posed human and object 3D model $(m^{i},m_{o}^{i})$, which are then sent to the cross-attention layers as guidance, down-sampled to match the resolution of each layer. Specifically, we create an input attention matrix $A\in\mathbb{R}^{{N_{i}}\times{N_{t}}}$ from the masks, applied to the cross-attention layers to encourage attention towards the intended region. We also modify the original procedure through the usage of a negative mask; specifically, we create a pseudo object segmentation map $m_{no}^{i}$ which, instead of using the intended hand, is using the opposite hand to grasp the 3D object model. This segmentation mask is then subtracted from the input attention matrix, disencouraging the generation in unintended locations.

### v3

@@ Abstract
Recent generative models can synthesize high-quality images, but they often fail to generate humans interacting with objects using their hands. This arises mostly from the model’s misunderstanding of such interactions and the hardships of synthesizing intricate regions of the body. In this paper, we propose GraspDiffusion, a novel generative method that creates realistic scenes of human-object interaction. Given a 3D object, GraspDiffusion constructs whole-body poses with control over the object’s location relative to the human body, which is achieved by separately leveraging the generative priors for body and hand poses, optimizing them into a joint grasping pose. This pose guides the image synthesis to correctly reflect the intended interaction, creating realistic and diverse human-object interaction scenes. We demonstrate that GraspDiffusion can successfully tackle the relatively uninvestigated problem of generating full-bodied human-object interactions while outperforming previous methods. Our project page is available at https://yj7082126.github.io/graspdiffusion/
Figure 1: Given an object mesh and its relative position, GraspDiffusion generates whole body grasping 3D poses, which is subsequently used as guidance for creating human-object interaction scenes. As shown, GraspDiffusion can synthesize images with valid human-object interactions for various types of objects. Note that the bottom-right sample (a green bag) was created from an object image, which was made into a 3D using TripoSR [TripoSR2024], further paving the way for various use cases.
11footnotetext: Work done while at Naver Webtoon.
@@ 1 Introduction
The recent advent of diffusion-based generative models [pmlr-v37-sohl-dickstein15, song2021scorebased, Ho2020DenoisingDP] has demonstrated significant success in producing high-quality visual content [Rombach2021HighResolutionIS, pmlr-v162-nichol22a, Ramesh2022Dalle2, saharia2022photorealistic, podell2024sdxl]. When trained on large datasets, these models can coherently synthesize images of various subjects corresponding to given textual/visual cues. However, despite their strong performance, generative models struggle to comprehend and visualize everyday hand-object interactions. This limitation hinders their broader adoption in generative model-based content creation pipelines.
@@
While several papers [lu2023handrefiner, wang2024rhandsrefiningmalformedhands, Pelykh2024GivingAH] have applied inpainting for hand region refinement, they only focus on situations where the hand is not interacting with other objects, making it impractical for most use cases. In order to generate realistic grasping hands for a given object, the models must understand the semantics and functionalities provided by the object–a concept well known as affordance [Gibson1979TheEA]. While Affordance Diffusion [Ye2023AffordanceDS] and HOIDiffusion [zhang2024hoidiffusion] creates an image of a single grasping hand using affordance, these methods only represent the explicit physical contact between a human hand and the object (devoid of any human identity), and fail to convey the spatial / orientational non-contact relationships, making them unsuitable for understanding human affordance. For instance, when using a cell phone, the relationships between the human’s face and torso with respect to the phone should be considered as part of interaction along with the hand touching the phone. This requirement necessitates the development of an pipeline that creates identifiable human-object images where the human’s body is visible, such that the implicit, non-contact relationships are well captured.
In this paper, we present GraspDiffusion, a novel method for generating interaction images with realistic hands and a clear human identity from a single 3D object input. Instead of relying on textual prompts, which tend to be ambiguous in describing complex interactions [Narasimhaswamy_2024_CVPR], we first utilize a diffusion model to generate full-bodied grasping poses [SMPL-X:2019, MANO:SIGGRAPHASIA:2017] conditioned on the object and its position. The generated 3D pose parameters are provided as conditions for the next stage to create high-quality image samples. Specifically, we train multiple conditional models [Zhang_2023_ICCV, Mou2023T2IAdapterLA] and a novel cross-attention modulation scheme [balaji2022eDiff-I, Couairon2023ZeroshotSL] to correctly convey the interaction context without harming the diversity of the generated samples. To overcome the lack of a large-scale image-3D pose paired dataset for human object interaction, we also propose a dataset annotation scheme to gather 3D annotations from previous image-based interaction datasets [Gupta2015VCOCO, Gkioxari2017DetectingAR, Chao2017HICODET, Li2019HAKEHA, Jin2020WholeBodyHP].
@@ 2 Related Work
Conditional Image Generation. To provide additional fine-grained, spatial conditions for diffusion models, ControlNet [Zhang_2023_ICCV] and T2I-Adapter [Mou2023T2IAdapterLA] proposed using image-level signals to control the generation process, which includes using 2D human keypoint skeletons [Cao2018OpenPoseRM, Yang2023EffectiveWP] for human pose guidance and depth maps for better depth perception. While the improvements in conditional image generation have been significant, synthesizing humans interacting with objects hasn’t reached the same level of improvement.
@@
Several papers [lu2023handrefiner, wang2024rhandsrefiningmalformedhands, Pelykh2024GivingAH] proposed to refine the hands of images generated using Controlnet-based inpainting, but does not count as a direct solution to human-object interaction. Others focused on identifying contacts for inpainting a new hand or object for a given scene [Ye2023AffordanceDS, Kulal2023PuttingPI, Min2024GenHeldGA, xue2024hoiswap], yet they are limited in camera views and require a prior scene context, making them infeasible in practical scenarios. HOIDiffusion [zhang2024hoidiffusion] also generates images from synthesized grasps, but is limited to hand-centric views. Compared to HanDiffuser [Narasimhaswamy_2024_CVPR], which applied the injection of hand embeddings during image generation to create realistic hands, our pipeline focuses more on the joint synthesis of hand and object, and uses spatial guidance to better direct the generation process towards a 3D scene context.
@@
Grasp Synthesis. Synthesizing a hand grasp consisting of a given object and a hand model is important in understanding human-object interaction, and is a widely studied task in robotics, graphics, and computer vision. While the focus in robotics is to make stable grasps for a given object in simulation / real life [Li2022GenDexGraspGD, Urain2022SE3DiffusionFieldsLS, Wang2022DexGraspNetAL, Geng2023UniDexGraspID, Weng2024DexDiffuserGD], in computer vision and graphics the focus is to make plausible grasps that are physically plausible, and generate grasps for either hands [Christen2021DGraspPP, Jiang2021GraspTTA, Karunratanakul2020GraspingFL, Zhou2022TOCHSO, Zheng2023CAMSCM, Christen2024DiffH2ODS, Liu2024GeneOHDT, Li2024ClickDiffCT] or a human [Tendulkar2022FLEXFG, Ghosh2022IMoSIF, Taheri2021GOALG4, Wu2021SAGASW, Braun2023PhysicallyPF, Zheng_2023_ICCV, Luo2024GraspingDiverse].
@@
Thanks to the advent of human-object interaction datasets [Bhatnagar2022BEHAVEDA, Brahmbhatt_2020_ECCV, Chao2021DexYCBAB, Fan2022ARCTICAD, Guzov2022VisuallyPH, Hampali2020honnotate, Kwon2021H2OTH, Liu2022HOI4DA4, Taheri2020GRABAD], many grasp synthesis methods achieved high performance in generating plausible grasps. Yet most existing datasets have issues with data scalability and variability, especially when it comes to color image-3D paired datasets. While hand-object interaction datasets like DexYCB [Chao2021DexYCBAB], ARCTIC [Fan2022ARCTICAD] and HOI4D [Liu2022HOI4DA4] consist of richly annotated image data, they are rather focused only on the hand and object recorded in an egocentric manner, devoid of any human identity and spatial / orientational non-contact relationships. Although BEHAVE [Bhatnagar2022BEHAVEDA] contains both RGB video sequences with 3D annotation, the image quality and motion sensors worn by the subject make it difficult to use as a realistic image dataset. To overcome this issue, we used traditional human-object interaction datasets [Gupta2015VCOCO, Gkioxari2017DetectingAR, Chao2017HICODET, Li2019HAKEHA, Jin2020WholeBodyHP] along with annotation tools [Kirillov2023SegmentA, liu2023grounding, li2021hybrik, li2023hybrikx, ke2023marigold] to create pseudo-3D annotations for the 2D image.
@@
Building our insight from similar approaches, [Tendulkar2022FLEXFG, Zheng_2023_ICCV], we focus on leveraging priors from a full-body pose model and a hand-grasping model. Compared to previous approaches [Taheri2021GOALG4, Ghosh2022IMoSIF, Wu2021SAGASW, Braun2023PhysicallyPF], instead of generating a motion sequence, we synthesize the pose parameters for the 3D parametric models [SMPL-X:2019, MANO:SIGGRAPHASIA:2017] using a diffusion model.
@@ 3 GraspDiffusion
Fig. 3 illustrates the proposed architecture. Starting with a 3D object mesh and its position within the human-centric coordinate system (originating at the pelvis joint), GraspDiffusion synthesizes realistic images portraying a human interacting with the object, with a significant portion of the human body visible to be considered ”full-body”. In the initial full-body grasping stage, we generate the pose parameters for the human body model [SMPL-X:2019] interacting with the 3D object mesh (Section. 3.2). In the scene generation stage, we extract geometric structures from the pose parameters to guide the generation of realistic images, leveraging a latent diffusion model [Rombach2021HighResolutionIS] along with spatial encoders and a cross-attention modulation scheme (Section. 3.3).
@@ 3.1 Preliminaries
Diffusion Models. Diffusion models [pmlr-v37-sohl-dickstein15, Ho2020DenoisingDP] are a group of generative models that interpret the data distribution $p(x)$ as a sequential transformation from a tractable prior distribution $p(x_{T})\sim\mathcal{N}(0,I)$. During training, the model uses a forward noise process $q(x_{t}|x_{t-1})$ that gradually adds a small amount of noise to a clean data sample $x_{0}$ towards $p(x_{T})$. At the same time, the model learns a backward noise process $p(x_{t-1}|x_{t})$ implemented as a neural network, which is trained to remove the noise from the before generating samples from $p(x_{T})$. For latent diffusion models [Rombach2021HighResolutionIS], the diffusion process is performed in the latent space of a trained autoencoder model [Kingma2013VAE], guided by a conditional text embedding derived from the CLIP [Radford2021CLIP] mechanism.
3D Parametric Models. For the hand model, we use the MANO [MANO:SIGGRAPHASIA:2017] differentiable model, in which we input the full finger articulated hand pose $\theta_{\text{h}}\in\mathbb{R}^{15\times 3}$, wrist translation $t_{\text{h}}\in\mathbb{R}^{3}$ and global orientation $R_{\text{h}}\in\mathbb{R}^{3}$ and get a 3D mesh $\mathcal{M}_{\text{h}}$ with vertices $\mathcal{V}_{\text{h}}$. For the full-body model, we use the SMPL-X [SMPL-X:2019] differentiable model, in which we input the full-body pose $\theta_{\text{b}}\in\mathbb{R}^{21\times 3}$, the full finger articulated hand pose $\theta_{\text{h}}\in\mathbb{R}^{15\times 3}$ for both hands, the root translation $t_{\text{b}}\in\mathbb{R}^{3}$ and global orientation $R_{\text{b}}\in\mathbb{R}^{3}$ and get a 3D mesh $\mathcal{M}_{\text{body}}$ with vertices $\mathcal{V}_{\text{b}}$.
Figure 4: Full-body grasping pipeline. We separately leverage a hand-grasping model [Taheri2020GRABAD] and a body-pose diffusion model, and perform a joint optimization into a full-bodied grasping pose.
@@ 3.2 Full-Body Grasping Pipeline
Building on prior approaches [Tendulkar2022FLEXFG, Zheng_2023_ICCV], we separately generate hand grasps and body poses in creating a whole-body grasping pose. Specifically, we take a 3D object mesh, its relative location to the human root, and the contacting hand orientation (left or right) as the input, to generate an SMPL-X mesh that grasps the given 3D object with a realistic body pose and hand-object contact.
The input object mesh is used to generate a plausible MANO [MANO:SIGGRAPHASIA:2017] hand grasp, for which we utilized GrabNet [Taheri2020GRABAD], a conditional variational autoencoder (cVAE) that produces hand grasps conditioned on the Basis Point Set (BPS) [Prokudin2019EfficientLO] of the given object. Separately trained for left and right-hand grasps, GrabNet generates MANO parameters $(\theta_{\text{h}},t_{\text{h}},R_{\text{h}})$ that grasps the given object, displaying accurate contact and high generalization for unseen objects.
The object’s relative location $t_{\text{obj}}\in\mathbb{R}^{3}$ and the hand orientation $c_{\text{left}},c_{\text{right}}\in\{0,1\}$ is then used to create a body pose that not only roughly positions its hand in the desired object location, but also reflects the appropriate implicit relationships required for a plausible grasping body pose [Taheri2021GOALG4, Kim2024BeyondTC]; whether the head is correctly oriented towards the object, the arms are correctly extended and the torso is leaning towards the object. To achieve this, we utilize a diffusion generative model trained to generate SMPL-X pose parameters $(\theta_{\text{body}},R_{\text{body}})$ conditioned on an object location and whether to use the right/left hand for contact. The loss is defined as
[…]
@@ 3.3 Scene Generation Pipeline
Given the 3D body pose pose, the scene generation pipeline extracts multiple spatial conditions as conditions to a pre-trained [Rombach2021HighResolutionIS] model to create consistent images of human-object interaction. Optionally, the pipeline can refine the image focused on the hand-object region to further adjust interaction and correct erroneous details.
@@
To precisely control the human-object image’s generation, we render three spatial conditions from the full-body grasping output. We first render the skeleton $(s^{i})$ of the SMPL-X body, consisting of body and hand joints, to ensure realistic human proportions within the generated image. We also use the joint depth map $(d^{i})$ from the SMPL-X and object model to provide depth information. Lastly, we render the occluded object with ambient lighting $(o^{i})$ to preserve its appearance while relighting it. To apply conditions, we chose the CoAdapter [Mou2023T2IAdapterLA] approach, which allows flexibility in handling multiple conditions. For each condition, we separately apply an adapter $\mathcal{F}_{\text{AD}}$ and perform a weighted sum to create feature $\mathbf{F}_{c}$, which can be written as
$\mathbf{F}_{c}=\sum_{k\in\{s,d,o\}}\omega_{k}\mathcal{F}_{\text{AD}}^{k}(k^{i}).\vskip-5.69046pt$ (3)
During training of the conditioning pipeline, we fix the parameters of the baseline U-Net model and only optimize the conditional adapters, reducing the risk of the model converging to the dataset’s style. This decision allows us to control the image style by applying LoRAs [hu2022lora] or using finetuned Stable Diffusion models during inference time, making it suitable for real-world application tasks requiring personalized image generation.
[…]
@@
To address the issue of erroneous interactions, in which interactions may occur from locations other than the intended area (examples on Fig. 8), we introduce a training-free guidance method motivated from prior zero-shot semantic image synthesis techniques  [balaji2022eDiff-I, Couairon2023ZeroshotSL]. We first render binary segmentation masks from the posed human and object 3D model $(m^{i},m_{o}^{i})$, which are then sent to the cross-attention layers as guidance, down-sampled to match the resolution of each layer. Specifically, we create an input attention matrix $A\in\mathbb{R}^{{N_{i}}\times{N_{t}}}$ from the masks, applied to the cross-attention layers to encourage attention towards the intended region. We also modify the original procedure through the usage of a negative mask; specifically, we create a pseudo object segmentation map $m_{no}^{i}$ which, instead of using the intended hand, is using the opposite hand to grasp the 3D object model. This segmentation mask is then subtracted from the input attention matrix, disencouraging the generation in unintended locations.